A Bibliography of Publications of Aaron Harwood

Aaron Harwood Griffith University School of Computers and Information Technology Australia Tel: ?n/a? FAX: ?n/a? E-mail: A.Harwood@cit.gu.edu.au (Internet) 25 November 2011 Version 1.01

Abstract

This bibliography records publications of Aaron Harwood.

Title word cross-reference

Υ [HS98a]. Basis [HS98a]. Class [HS98a]. Cost [HS98b]. Electrical [HS97]. Expanded [HS98a]. Fat [HS98c]. Fat-tree [HS98c]. Fundamental [HS97]. Highly [HS98b]. Hybrid [HS98c]. Interconnection [HS98a]. Low [HS98b].

Network [HS98b]. Networks [HS98a].

Quantum [HS97].

Scalable [HS98b]. Scheduling [HS97].

Theory [HS97]. Time [HS97]. tree [HS98c]. Uniprocessor [HS97]. Using [HS97]. Varying [HS97].

References

Harwood:1997:UFE

[HS97] Aaron Harwood and Hong Shen. Using fundamental electrical theory for varying time quantum uniprocessor scheduling. In Region Ten Conference of the IEEE Speech and Image Technologies for Computing and Telecommunications. Griffith University, School of Computers and Information Technology and School of Microelectronic Engineering, Brisbane, Australia, 1997.

Harwood:1998:CIN

[HS98a] Aaron Harwood and Hong Shen. A class of interconnection networks on the basis of expanded Υ networks. In Proceedings of International Conference on Parallel and Distributed Processing and Applications. Griffith University, School of Computers and Information Technology and School of Microelectronic Engineering, Brisbane, Australia, 1998.

Harwood:1998:HSL

[HS98b] Aaron Harwood and Hong Shen. A highly scalable and low cost interconnection network. In 2^{nd} International Conference on Parallel and Distributed Computing and Networks. Griffith University, School of Computers and Information Technology and School of Microelectronic Engineering, Brisbane, Australia, 1998.

Harwood:1998:LCH

[HS98c] Aaron Harwood and Hong Shen.

A low cost hybrid fat-tree interconnection network. In Proceedings of International Conference on Parallel and Distributed Processing and Applications. Griffith University, School of Computers and Information Technology and School of Microelectronic Engineering, Brisbane, Australia, 1998.